# Mode analysis of Nambu-Goldstone modes in U(1) charged first-order relativistic viscous hydrodynamics

### v2

@@ 6 Conclusion
While some of the results recovered in this work were known in previous literature based on the traditional mode analysis of hydrodynamical perturbations, our approach is novel. For instance, the stability of long-wavelength hydrodynamical modes is widely recognized. However, to the best of our knowledge, we have clearly addressed the hydrodynamical frame invariance for a generic $U(1)$ fluid with a proper gradient expansion for the first time. Additionally, we found that higher-order gradients enter the equation of motion when solving the constraint equations, even if the constitutive relations are first-order. Furthermore, we directly applied the results of Ref. Heller:2022ejw to derive the bound on the adiabatic sound speed, and we explicitly demonstrated that the assumption in Ref. Heller:2022ejw can be satisfied in a frame-invariant manner under arbitrary Lorentz boosts.
For any thermal UV theory underlying hydrodynamics with discrete symmetries, the conjugation, unitarity, and KMS conditions must be satisfied. Hence, the additional requirement for stability reduces to $\rho+P\geq 0$, known as the null energy condition in a Minkowski background.
A key advantage of our construction is the incorporation of techniques from nonlinear realization theory. By construction, our hydrodynamical action is invariant under Lorentz boosts, so the stability analysis is valid in any Lorentz frame. In our framework, the Lorentz boost transformation is nonlinearly realized, ensuring that the adiabatic sound speed remains unchanged. This implies that the intrinsic sound speed is always defined with respect to the rest frame of the fluid. The relative sound speed due to fluid motion should be distinct from the adiabatic sound speed. To summarize, we assert that a generic thermal $U(1)$-charged first-order hydrodynamics in a Minkowski background is stable in the hydrodynamical regime, provided that the null energy condition is satisfied.
Finally, we commented on the causal structure for the hydrodynamical modes. Without dissipation, the theory is identical to free scalar field theory in a Minkowski background, and thus the retarded Green function straightforwardly defines causality with respect to the sound cone. Propagation of the wavefront due to diffusion violates the causality structure for the sound cone. However, the wavefront propagates ahead due to the diffusion of fast modes rather than the sound wave. We claimed that the front velocity must be given by a composition of the intrinsic sound velocity $c_{s/n}\leq 1$ and the maximum velocity of the fast mode $c_{\rm fast}\leq 1$, which is constrained by the speed of light within special relativity. Hence, while the retarded correlation may extend beyond the sound cone, it should never exceed the light cone. Integrating the fast mode’s free-streaming into hydrodynamics will provide a consistent solution for the causality issue in first order hydrodynamics.
The nonlinear evolution of NG modes can be studied by expanding the action to higher orders in $\partial\pi$. Since the Green functions are hydrodynamical frame-invariant, the loop corrections will also be hydrodynamical frame-invariant if the interactions are similarly hydrodynamical frame-invariant, which we will discuss in future projects.
We found the stability of the NG modes in the hydrodynamical regime frame independent. If the dynamics of perturbations depend on the choice of the hydrodynamical frame, this may signal a breakdown of the hydrodynamical approximation. Hence, investigation of hydrodynamical perturbation could provide a consistency check in e.g., hydrodynamical simulations. In this paper, we applied the principle of hydrodynamical frame independence to guide our mode analysis, and as far as we are concerned with the hydrodynamical scale, it was successful. Consequently, we took a different approach to first-order hydrodynamics than that proposed in the BDNK theory Bemfica:2017wps ; Kovtun:2019hdm ; Bemfica:2019knx ; Bemfica:2020zjp , where hydrodynamical frame dependence plays a central role. Further exploration of the relationship between these approaches will be important.
Finally, we comment on several applications of the effective field theory of dissipative fluids in its present form. Our effective action closely resembles that of, for instance, the effective field theory of inflation Cheung:2007st , and its extensions to open systems Hongo:2018ant ; Salcedo:2024smn . In the context of inflation, time diffeomorphism symmetry broken by a background solution of the inflaton field is nonlinearly realized. In our case, all diffeomorphism symmetries are nonlinearly realized, with two components reduced to vector modes and one to longitudinal mode. The temporal mode is eliminated by solving the constraint equations. One could eliminate the longitudinal mode and obtain an action for the temporal NG modes. However, as we discussed in Sec. 4.3, additional fluid spacetime symmetries and first-order constitutive relations arise in the hydrodynamical setup. As a result, we argued that gapped modes do not appear, contrary to Ref. Hongo:2018ant .
@@
Extending our formulation to cosmological backgrounds, such as Schwarzschild or Friedmann-Lemaître-Robertson-Walker spacetimes, and coupling to dynamical freedom in the metric is straightforward. This approach could be particularly useful for an action-based formulation of the effective field theory of large-scale structure Carrasco:2012cv ; Porto:2013qua , including characteristic diffusion scales. Also the effective theory of sound modes will be useful for describing warm inflation Berera:1995wh ; Berera:1995ie ; Berera:2023liv .
Acknowledgements.
@@
Appendix A Useful equations
A.1 Thermodynamical relations
@@
    $\rho_{\beta}-\mu n_{\beta}=\frac{s_{\beta}}{\beta},~{}\rho_{\mu}-\mu n_{\mu}=\frac{s_{\mu}}{\beta}.$ (271)
@@
    $P_{\beta}$ $=-\frac{s}{\beta^{2}},~{}P_{\mu}=n,$ (272)
@@
    $-\frac{s_{\mu}}{\beta^{2}}=n_{\beta}.$ (273)
A.2 KMS condition
@@
$\rho_{(1)}$ $=\beta^{-1}\epsilon u^{\rho}\nabla_{\rho}\beta-\lambda\nabla_{\rho}u^{\rho}-\nu_{1}u^{\rho}\nabla_{\rho}(\beta\mu),$ (274)
$P_{(1)}$ $=\beta^{-1}\lambda u^{\rho}\nabla_{\rho}\beta-\zeta\nabla_{\rho}u^{\rho}-\nu_{2}u^{\rho}\nabla_{\rho}(\beta\mu),$ (275)
$n_{(1)}$ $=\nu_{1}u^{\rho}\nabla_{\rho}\beta-\nu_{2}\beta\nabla_{\rho}u^{\rho}-\nu_{3}\beta u^{\rho}\nabla_{\rho}(\beta\mu),$ (276)
$q_{(1)\mu}$ $=\kappa(\beta^{-1}\nabla_{\mu}\beta-u^{\rho}\nabla_{\rho}u_{\mu})-\tau(\nabla_{\mu}(\beta\mu)+\beta u^{\rho}F_{\rho\mu}),$ (277)
$j_{(1)\mu}$ $=\tau(\nabla_{\mu}\beta-\beta u^{\rho}\nabla_{\rho}u_{\mu})-\chi(\beta\nabla_{\rho}(\beta\mu)+\beta^{2}u^{\rho}F_{\rho\mu}),$ (278)
$\begin{split}\beta\hbar W_{(0)}^{\mu\nu\rho\sigma}&=\epsilon u^{\mu}u^{\nu}u^{\rho}u^{\sigma}+\lambda(\gamma^{\mu\nu}u^{\rho}u^{\sigma}+\gamma^{\rho\sigma}u^{\mu}u^{\nu})\\ &+\kappa(\gamma^{\mu\rho}u^{\nu}u^{\sigma}+\gamma^{\nu\rho}u^{\mu}u^{\sigma}+\gamma^{\mu\sigma}u^{\nu}u^{\rho}+\gamma^{\nu\sigma}u^{\mu}u^{\rho})\\ &+\zeta\gamma^{\mu\nu}\gamma^{\rho\sigma}+\eta\left(\gamma^{\mu\rho}\gamma^{\nu\sigma}+\gamma^{\nu\rho}\gamma^{\mu\sigma}-\frac{2}{3}\gamma^{\mu\nu}\gamma^{\rho\sigma}\right),\end{split}$ (279)
$\begin{split}\beta\hbar X_{(0)}^{\mu\nu\rho}&=\nu_{1}u^{\mu}u^{\nu}u^{\rho}+\nu_{2}\gamma^{\mu\nu}u^{\rho}+\tau(\gamma^{\mu\rho}u^{\nu}+\gamma^{\nu\rho}u^{\mu}),\end{split}$ (280)
$\begin{split}\beta\hbar Y_{(0)}^{\mu\rho}&=\nu_{3}u^{\mu}u^{\rho}+\chi\gamma^{\mu\rho}.\end{split}$ (281)
A.3 Adiabatic sound velocity
@@
$=\left(\frac{\partial P}{\partial\rho}\right)_{s/n}\left(\rho_{\mu}d\mu+\rho_{\beta}d\beta\right)+\frac{s}{n}\left(\frac{\partial P}{\partial(s/n)}\right)_{\rho}\left(\frac{s_{\mu}d\mu+s_{\beta}d\beta}{s}-\frac{n_{\mu}d\mu+n_{\beta}d\beta}{n}\right).$ (282)
@@
$\left(\frac{\partial P}{\partial\rho}\right)_{s/n}$ $=\frac{n(P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu})+(n_{\mu}P_{\beta}-n_{\beta}P_{\mu})(\rho+P)}{(\rho+P)(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}=c_{s/n}^{2}.$ (283)
@@
$\left(\frac{\partial P}{\partial\rho}\right)_{s}$ $=\frac{\mu(n_{\mu}P_{\beta}-n_{\beta}P_{\mu})+P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu}}{\mu(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}=c^{2}_{s},$ (284)
$\left(\frac{\partial P}{\partial\rho}\right)_{n}$ $=\frac{n_{\mu}P_{\beta}-n_{\beta}P_{\mu}}{n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu}}=c_{n}^{2}.$ (285)
@@
$c_{s/n}^{2}=\frac{nc_{s}^{2}+sc_{n}^{2}}{s+\beta\mu n}=\frac{nc_{s}^{2}+sc_{n}^{2}}{\rho+P}.$ (286)
@@
$\left(\frac{\partial P}{\partial n}\right)_{\rho}$ $=\frac{P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu}}{n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu}},$ (287)
$\left(\frac{\partial P}{\partial s}\right)_{\rho}$ $=-\frac{P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu}}{\beta\mu(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}.$ (288)
Appendix B 1D Brownian motion
@@
$m\ddot{X}+m\gamma\dot{X}+m\omega^{2}_{0}X=\xi,~{}\langle\xi(t)\xi(t^{\prime})\rangle=\frac{2m\gamma}{\beta}\delta(t-t^{\prime}).$ (289)
@@
$\ddot{G}+\gamma\dot{G}+\omega^{2}_{0}G=\delta(t-t^{\prime}),$ (290)
@@
Evaluating Eq. (291) by using Eq. (297), the Keldysh propagator is found as
@@
$G(t-t^{\prime})=\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}G(\omega)$ (294)
@@
$-\omega^{2}G(\omega)-i\omega\gamma G(\omega)+\omega^{2}_{0}G(\omega)=1,$ (295)
@@
$G(\omega)=\frac{1}{\omega_{0}^{2}-\omega^{2}-i\omega\gamma}.$ (296)
@@
$G(t)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{e^{-i\omega t}}{\omega_{0}^{2}-\omega^{2}-i\omega\gamma}=\Theta(t)\frac{e^{-\frac{\gamma t}{2}}}{\omega_{0}}\frac{\sin\left(\omega_{0}t\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}\right)}{\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}}.$ (297)
Finally, let us evaluate Eqs. (293) by using (297). We are particularly interested in the case where $(t+t_{1})/2\to\infty$ while $t-t_{1}$ is finite: sufficient time has passed since the initial time of the thermal state, but consider the correlation for a finite time interval. In this limit, the analytic expression simplifies, and in the small $\gamma$ limit, we find the following asymptotic form:
$\langle X(t)X(t_{1})\rangle$ $\simeq\frac{e^{-\frac{1}{2}\gamma|t-t_{1}|}}{m\beta\omega_{0}^{2}}\left(\cos\left(\omega_{0}|t-t_{1}|\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}\right)+\frac{\gamma}{2\omega_{0}}\frac{\sin\left(\omega_{0}|t-t_{1}|\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}\right)}{\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}}\right).$ (298)
@@
Let us find the same expression from the path integral. The MSR action for Eq. (289) is written as
@@
$Y=\binom{y}{{\bar{y}}},~{}J=\binom{\bar{j}}{j},~{}B=\begin{pmatrix}0&\omega^{2}-\omega^{2}_{0}-i\omega\gamma\\ \omega^{2}-\omega^{2}_{0}+i\omega\gamma&\frac{2i\gamma}{\beta}\end{pmatrix}$ (303)
@@
$iS$ $=\frac{i}{2}\int\frac{d\omega}{2\pi}Y^{\dagger}BY+J^{\dagger}Y+Y^{\dagger}J,$ (304)
@@
$Y\to Y-B^{-1}J,~{}iB^{-1}=\begin{pmatrix}\frac{2\gamma}{\beta\left(\gamma^{2}\omega^{2}+\left(\omega^{2}-\omega_{0}^{2}\right)^{2}\right)}&\frac{1}{\gamma\omega-i\left(\omega^{2}-\omega_{0}^{2}\right)}\\ -\frac{1}{\gamma\omega+i\left(\omega^{2}-\omega_{0}^{2}\right)}&0\end{pmatrix}$ (305)
@@
$=$ $Y^{\dagger}BY-J^{\dagger}(B^{-1})^{\dagger}BY+J^{\dagger}Y-J^{\dagger}B^{-1}J.$ (306)
For a Hermitian $B$, we get $(B^{-1})^{\dagger}B=1$, and then $J^{\dagger}Y$ is subtracted. In our case, $B$ is not Hermitian: $(B^{-1})^{\dagger}B\neq 1$. We are not able to integrate $y$ independently from $j$ due to this nonvanishing cross-term. In fact, Eq. (305) is not a unique way to find the bilinear form of $J$. Instead, one may consider $Y^{\dagger}\to Y^{\dagger}-Y^{\dagger}B^{-1}$. Then we find $Y^{\dagger}J-Y^{\dagger}B(B^{-1})^{\dagger}J$ remains. These terms are $\gamma/\beta$ multiplied by the retarded and advanced propagators, respectively. Then, we may split the path integral for $t<0$ and $t>0$ and eliminate these cross terms. Finally, we find the Green functions found in Eqs. (297) and (298) after rescaling by $m$:
$\int\frac{d\omega}{2\pi}e^{-i\omega t}iB^{-1}=\begin{pmatrix}G_{K}&G_{R}\\ G_{A}&0\end{pmatrix},$ (307)

### v3

@@ 7 Conclusions
Some of the results obtained in this work were known from earlier analyses of hydrodynamical perturbations using traditional methods. However, our approach differs in several important aspects. For instance, although the stability of long-wavelength hydrodynamical modes is widely recognized, to the best of our knowledge, this is the first work to explicitly establish hydrodynamic-frame invariance for a generic $U(1)$ fluid using a consistent gradient expansion. We also found that higher-order gradients can enter the equations of motion through the solution of constraint equations, even if the constitutive relations are first order.
A key advantage of our framework is the use of techniques from nonlinear realization theory. The hydrodynamical action is constructed to be invariant under nonlinearly realized Lorentz boosts, ensuring that the stability analysis is valid in any Lorentz frame. In our formulation, Lorentz boosts are realized nonlinearly, which ensures that the adiabatic sound speed remains unchanged. This means that the intrinsic sound speed is always defined in the rest frame of the fluid. The apparent sound speed arising from fluid motion should be distinguished from the adiabatic sound speed. One might expect that sound waves before and after a Lorentz transformation are related through a transformation of the wavevector, $k_{\mu}\to k^{\prime}_{\mu}=\Lambda_{\mu}{}^{\nu}k_{\nu}$. This expectation holds for linearly realized representations such as photons. In contrast, sound waves are nonlinear realizations, and identifying them through linear transformations mixes variables that are not physically equivalent from a covariant perspective. This is the reason why mode analysis in the traditional approach depends on the choice of Lorentz frames.
A natural direction for future work is to study the nonlinear evolution of Nambu–Goldstone modes. Since the first-order theory is now well-defined within the hydrodynamical regime, an extension to loop-level analysis becomes straightforward. Because the Green’s functions are hydrodynamic-frame invariant, loop corrections will also be invariant if the interactions are constructed to respect the same invariance. This aspect will be discussed in future work.
In this paper, we applied the principle of hydrodynamic-frame independence to guide our mode analysis. Within the hydrodynamical regime, this approach proved successful. As a result, our perspective on first-order hydrodynamics differs from that of the BDNK theory [36, 23, 37, 34], where frame dependence plays a central role. A more detailed comparison between these approaches will be important in future studies.
@@
Finally, we comment on possible applications of our effective field theory of dissipative fluids. Our effective action shares structural similarities with that of the effective field theory of inflation [38], and with its extensions to open systems [30, 39]. The effective theory of sound modes presented here may also be useful for describing warm inflation [40, 41, 42]. It is straightforward to generalize our framework to cosmological backgrounds, such as Schwarzschild or Friedmann–Lemaître–Robertson–Walker spacetimes, and to couple it to the dynamical metric degrees of freedom. Our approach may be particularly suited to an action-based formulation of the effective field theory of large-scale structure [43, 44], since it does not rely on thermal equilibrium assumptions at the outset.
Acknowledgments
@@
Appendix A Notations
As relativistic hydrodynamics is interdisciplinary and uses many symbols to describe in a general setup, depending on the contexts and authors, equivalent expressions are expressed in multiple ways. This appendix provides a table of translation of hydrodynamical coefficients between our work and a seminal reference [23]. We summarize the correspondence of variables and transport coefficients in Table 1. This allows for a direct comparison of constitutive relations and facilitates the identification of frame-invariant quantities. Here, the frame invariants such as $f_{i}$ and $\ell_{i}$ are defined in Ref. [23] as combinations of transport coefficients that are invariant under first-order field redefinitions of hydrodynamical variables. See Eq. (2.10) of Ref. [23]. Note that the external vector field is not introduced in Ref. [23].
Table 1: Correspondence of first-order hydrodynamical coefficients. Frame invariants (on-shell) means combinations relevant for observable response functions after eliminating redundant degrees of freedom using the equations of motion.
Physical quantity Our notation Ref. [23]
Temperature $\beta^{-1}$ $T$
Chemical potential $\mu$ $\mu$
4-velocity $u^{\mu}$ $u^{\mu}$
Adiabatic sound speed $c_{s},c_{s/n}$ $v_{s}$
Spacetime metric $g_{\mu\nu}$ $\eta_{\mu\nu}$ (Minkowski)
Induced metric $\gamma_{\mu\nu}$ $\Delta_{\mu\nu}$
Vector field $A_{\mu}$ none
Energy density $(\rho,\epsilon,\lambda_{1},\nu_{4})$ $(\epsilon,-\varepsilon_{1},-\varepsilon_{2},-\varepsilon_{3})$
Pressure $(P,\lambda_{2},\zeta,\nu_{5})$ $(p,-\pi_{1},-\pi_{2},-\pi_{3})$
Charge density $(n,\nu_{1},\nu_{2},\nu_{3})$ $(n,-\nu_{1},-\nu_{2},-\nu_{3})$
Momentum density $(\kappa_{1},\kappa_{2},\tau_{3},\tau_{4})$ $(-\theta_{2},-\theta_{1},-\theta_{3},0)$
Charge flux $(\tau_{1},\tau_{2},\chi_{1},\chi_{2})$ $(-\gamma_{2}T,-\gamma_{1}T,-\gamma_{3}T,0)$
Shear viscosity $\eta$ $-\eta$
Frame invariants $\left(\tilde{\lambda},\tilde{\zeta},\tilde{\nu},\tilde{\tau}_{1},\tilde{\tau}_{2},\tilde{\chi}_{1},\tilde{\chi}_{2}\right)$ $(-f_{1},-f_{2},-f_{3},-\ell_{2}T,-\ell_{1}T,-\ell_{3}T,0)$
Frame invariants (on-shell) $\left({\bm{\zeta}},{\bm{\chi}}_{1},{\bm{\chi}}_{2},{\bm{\tau}}\right)$ $(\zeta,-T^{2}\sigma,{\rm none},T^{2}\chi_{\rm T})$
Appendix B Useful equations
B.1 Thermodynamical relations
@@
    $\rho_{\beta}-\mu n_{\beta}=\frac{s_{\beta}}{\beta},~\rho_{\mu}-\mu n_{\mu}=\frac{s_{\mu}}{\beta}.$ (B.1)
@@
    $P_{\beta}$ $=-\frac{s}{\beta^{2}},~P_{\mu}=n,$ (B.2)
@@
    $-\frac{s_{\mu}}{\beta^{2}}=n_{\beta}.$ (B.3)
B.2 KMS condition
@@
$\rho_{(1)}$ $=\beta^{-1}\epsilon u^{\rho}\nabla_{\rho}\beta-\lambda\nabla_{\rho}u^{\rho}-\nu_{1}u^{\rho}\nabla_{\rho}(\beta\mu),$ (B.4)
$P_{(1)}$ $=\beta^{-1}\lambda u^{\rho}\nabla_{\rho}\beta-\zeta\nabla_{\rho}u^{\rho}-\nu_{2}u^{\rho}\nabla_{\rho}(\beta\mu),$ (B.5)
$n_{(1)}$ $=\nu_{1}u^{\rho}\nabla_{\rho}\beta-\nu_{2}\beta\nabla_{\rho}u^{\rho}-\nu_{3}\beta u^{\rho}\nabla_{\rho}(\beta\mu),$ (B.6)
$q_{(1)\mu}$ $=\kappa(\beta^{-1}\nabla_{\mu}\beta-u^{\rho}\nabla_{\rho}u_{\mu})-\tau(\nabla_{\mu}(\beta\mu)+\beta u^{\rho}F_{\rho\mu}),$ (B.7)
$j_{(1)\mu}$ $=\tau(\nabla_{\mu}\beta-\beta u^{\rho}\nabla_{\rho}u_{\mu})-\chi(\beta\nabla_{\rho}(\beta\mu)+\beta^{2}u^{\rho}F_{\rho\mu}),$ (B.8)
$\begin{split}\beta\hbar W_{(0)}^{\mu\nu\rho\sigma}&=\epsilon u^{\mu}u^{\nu}u^{\rho}u^{\sigma}+\lambda(\gamma^{\mu\nu}u^{\rho}u^{\sigma}+\gamma^{\rho\sigma}u^{\mu}u^{\nu})\\ &+\kappa(\gamma^{\mu\rho}u^{\nu}u^{\sigma}+\gamma^{\nu\rho}u^{\mu}u^{\sigma}+\gamma^{\mu\sigma}u^{\nu}u^{\rho}+\gamma^{\nu\sigma}u^{\mu}u^{\rho})\\ &+\zeta\gamma^{\mu\nu}\gamma^{\rho\sigma}+\eta\left(\gamma^{\mu\rho}\gamma^{\nu\sigma}+\gamma^{\nu\rho}\gamma^{\mu\sigma}-\frac{2}{3}\gamma^{\mu\nu}\gamma^{\rho\sigma}\right),\end{split}$ (B.9)
$\begin{split}\beta\hbar X_{(0)}^{\mu\nu\rho}&=\nu_{1}u^{\mu}u^{\nu}u^{\rho}+\nu_{2}\gamma^{\mu\nu}u^{\rho}+\tau(\gamma^{\mu\rho}u^{\nu}+\gamma^{\nu\rho}u^{\mu}),\end{split}$ (B.10)
$\begin{split}\beta\hbar Y_{(0)}^{\mu\rho}&=\nu_{3}u^{\mu}u^{\rho}+\chi\gamma^{\mu\rho}.\end{split}$ (B.11)
B.3 Adiabatic sound velocity
@@
$=\left(\frac{\partial P}{\partial\rho}\right)_{s/n}\left(\rho_{\mu}d\mu+\rho_{\beta}d\beta\right)+\frac{s}{n}\left(\frac{\partial P}{\partial(s/n)}\right)_{\rho}\left(\frac{s_{\mu}d\mu+s_{\beta}d\beta}{s}-\frac{n_{\mu}d\mu+n_{\beta}d\beta}{n}\right).$ (B.12)
@@
$\left(\frac{\partial P}{\partial\rho}\right)_{s/n}$ $=\frac{n(P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu})+(n_{\mu}P_{\beta}-n_{\beta}P_{\mu})(\rho+P)}{(\rho+P)(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}=c_{s/n}^{2}.$ (B.13)
@@
$\left(\frac{\partial P}{\partial\rho}\right)_{s}$ $=\frac{\mu(n_{\mu}P_{\beta}-n_{\beta}P_{\mu})+P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu}}{\mu(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}=c^{2}_{s},$ (B.14)
$\left(\frac{\partial P}{\partial\rho}\right)_{n}$ $=\frac{n_{\mu}P_{\beta}-n_{\beta}P_{\mu}}{n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu}}=c_{n}^{2}.$ (B.15)
@@
$c_{s/n}^{2}=\frac{nc_{s}^{2}+sc_{n}^{2}}{s+\beta\mu n}=\frac{nc_{s}^{2}+sc_{n}^{2}}{\rho+P}.$ (B.16)
@@
$\left(\frac{\partial P}{\partial n}\right)_{\rho}$ $=\frac{P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu}}{n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu}},$ (B.17)
$\left(\frac{\partial P}{\partial s}\right)_{\rho}$ $=-\frac{P_{\mu}\rho_{\beta}-P_{\beta}\rho_{\mu}}{\beta\mu(n_{\mu}\rho_{\beta}-n_{\beta}\rho_{\mu})}.$ (B.18)
Appendix C 1D Brownian motion
@@
$m\ddot{X}+m\gamma\dot{X}+m\omega^{2}_{0}X=\xi,~\langle\xi(t)\xi(t^{\prime})\rangle=\frac{2m\gamma}{\beta}\delta(t-t^{\prime}).$ (C.1)
@@
$\ddot{G}+\gamma\dot{G}+\omega^{2}_{0}G=\delta(t-t^{\prime}),$ (C.2)
@@
The Keldysh propagator is found as
@@
$G(t-t^{\prime})=\int\frac{d\omega}{2\pi}e^{-i\omega(t-t^{\prime})}G(\omega)$ (C.6)
@@
$-\omega^{2}G(\omega)-i\omega\gamma G(\omega)+\omega^{2}_{0}G(\omega)=1,$ (C.7)
@@
$G(\omega)=\frac{1}{\omega_{0}^{2}-\omega^{2}-i\omega\gamma}.$ (C.8)
@@
$G(t)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{e^{-i\omega t}}{\omega_{0}^{2}-\omega^{2}-i\omega\gamma}=\Theta(t)\frac{e^{-\frac{\gamma t}{2}}}{\omega_{0}}\frac{\sin\left(\omega_{0}t\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}\right)}{\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}}.$ (C.9)
Finally, let us evaluate Eqs. (C.5) by using (C.9). We are particularly interested in the case where $(t+t_{1})/2\to\infty$ while $t-t_{1}$ is finite: sufficient time has passed since the initial time of the thermal state, but consider the correlation for a finite time interval. In this limit, the analytic expression simplifies, and in the small $\gamma$ limit, we find the following asymptotic form:
$\langle X(t)X(t_{1})\rangle$ $\simeq\frac{e^{-\frac{1}{2}\gamma|t-t_{1}|}}{m\beta\omega_{0}^{2}}\left(\cos\left(\omega_{0}|t-t_{1}|\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}\right)+\frac{\gamma}{2\omega_{0}}\frac{\sin\left(\omega_{0}|t-t_{1}|\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}\right)}{\sqrt{1-\frac{\gamma^{2}}{4\omega_{0}^{2}}}}\right).$ (C.10)
@@
Let us find the same expression from the path integral. The MSR action for Eq. (C.1) is written as
@@
$Y=\binom{y}{{\bar{y}}},~J=\binom{\bar{j}}{j},~B=\begin{pmatrix}0&\omega^{2}-\omega^{2}_{0}-i\omega\gamma\\ \omega^{2}-\omega^{2}_{0}+i\omega\gamma&\frac{2i\gamma}{\beta}\end{pmatrix}$ (C.15)
@@
$iS$ $=\frac{i}{2}\int\frac{d\omega}{2\pi}Y^{\dagger}BY+J^{\dagger}Y+Y^{\dagger}J,$ (C.16)
@@
$Y\to Y-B^{-1}J,~iB^{-1}=\begin{pmatrix}\frac{2\gamma}{\beta\left(\gamma^{2}\omega^{2}+\left(\omega^{2}-\omega_{0}^{2}\right)^{2}\right)}&\frac{1}{\gamma\omega-i\left(\omega^{2}-\omega_{0}^{2}\right)}\\ -\frac{1}{\gamma\omega+i\left(\omega^{2}-\omega_{0}^{2}\right)}&0\end{pmatrix}$ (C.17)
@@
$=$ $Y^{\dagger}BY-J^{\dagger}(B^{-1})^{\dagger}BY+J^{\dagger}Y-J^{\dagger}B^{-1}J.$ (C.18)
For a Hermitian $B$, we get $(B^{-1})^{\dagger}B=1$, and then $J^{\dagger}Y$ is subtracted. In our case, $B$ is not Hermitian: $(B^{-1})^{\dagger}B\neq 1$. We are not able to integrate $y$ independently from $j$ due to this nonvanishing cross-term. In fact, Eq. (C.17) is not a unique way to find the bilinear form of $J$. Instead, one may consider $Y^{\dagger}\to Y^{\dagger}-Y^{\dagger}B^{-1}$. Then we find $Y^{\dagger}J-Y^{\dagger}B(B^{-1})^{\dagger}J$ remains. These terms are $\gamma/\beta$ multiplied by the retarded and advanced propagators, respectively. Then, we may split the path integral for $t<0$ and $t>0$ and eliminate these cross terms. Finally, we find the Green functions found in Eqs. (C.9) and (C.10) after rescaling by $m$:
$\int\frac{d\omega}{2\pi}e^{-i\omega t}iB^{-1}=\begin{pmatrix}G_{K}&G_{R}\\ G_{A}&0\end{pmatrix},$ (C.19)